# Recurrence Patterns Correlation

### v1

@@ Abstract
Recurrence plots (RPs) are powerful tools for visualizing time series dynamics; however, traditional Recurrence Quantification Analysis (RQA) often relies on global metrics, such as line counting, that can overlook system-specific, localized structures. To address this, we introduce Recurrence Pattern Correlation (RPC), a quantifier inspired by spatial statistics that bridges the gap between qualitative RP inspection and quantitative analysis. RPC is designed to measure the correlation degree of an RP to patterns of arbitrary shape and scale. By choosing patterns with specific time lags, we visualize the unstable manifolds of periodic orbits within the Logistic map bifurcation diagram, dissect the mixed phase space of the Standard map, and track the unstable periodic orbits of the Lorenz ’63 system’s 3-dimensional phase space. This framework reveals how long-range correlations in recurrence patterns encode the underlying properties of nonlinear dynamics and provides a more flexible tool to analyze pattern formation in RPs.
@@ II Methods
is the average over all RP entries, known as the recurrence rate. The weights $\text{{w}}_{\Delta i,\Delta j}$ are defined over time lags $\Delta i=i^{\prime}-i$ and $\Delta j=j^{\prime}-j$, so the correlation structure is translation invariant over the entire RP, making RPC a global measure of such generic motif. Moreover, the quantifier is normalized given the factor
@@
Assuming the relatedness of time series values is local, we consider lags restricted to a time window outside which $\text{{w}}_{\Delta i,\Delta j}$ is null, so the second summation only encompasses non-null weight indexes. Otherwise, the method could be computationally unfeasible for large $\text{{w}}_{\Delta i,\Delta j}$.
@@
In this context, we aim to develop a quantifier that captures not only global recurrence patterns (RPC) but also those localized to individual values in the time series, and so in the state space. The local version can be conceptualized as applying the quantifier to individual columns of the RP, each associated with a specific time index $i$ in Eq. 2, and is defined as follows:
@@
As a result, the global RPC is not simply the average of the local values. This distinction makes the global definition better suited for identifying properties shared across the entire time series. While the local RPC captures context-sensitive features by adapting to the local recurrence rate.
@@ III Results and Discussion
For one-dimensional time series, we applied the PECUZAL algorithm [25] for phase space reconstruction. Which optimizes embedding parameters using non-uniform delays and a cost function tailored to the embedding dimension, without relying on a threshold parameter.
@@
The simplest correlated patterns arise from combinations of nearby recurrences, defined by the weight matrix $\text{{w}}_{\Delta i,\Delta j}=1$ only when $\Delta i=\pm 1$ or $\Delta j=\pm 1$. This corresponds to single time-lag recurrences aligned with vertical, diagonal, or anti-diagonal directions. Such a definition is commonly used in spatial correlation analysis, where long-range temporal causality is disregarded in favor of identifying nearby values in spatial clusters. We apply this simple weight matrix to representative dynamical systems (Fig. 3), revealing distinct recurrence fingerprints that differentiate stochastic from deterministic dynamics.
@@
Figure 3: Recurrence Patterns Correlation (RPC) for paradigmatic dynamics and representative time lags, considering a $1\%$ recurrence rate across all recurrence plots. In this scale, the recurring condition is highlighted over nonrecurring patterns. Each bar represents a specific weight matrix $\text{{w}}_{\Delta i,\Delta j}$, corresponding to: sides ($\text{{w}}_{\Delta i,\Delta j}=\delta_{\Delta i,0}\delta_{\Delta j,\pm 1}+\delta_{\Delta i,\pm 1}\delta_{\Delta j,0}$), diagonals ($\text{{w}}_{\Delta i,\Delta j}=\delta_{\Delta i,\pm 1}\delta_{\Delta j,\pm 1}$), and anti-diagonals ($\text{{w}}_{\Delta i,\Delta j}=\delta_{\Delta i,\pm 1}\delta_{\Delta j,\mp 1}$).
@@
Diagonal Motif: The correlation for the diagonal motif ($\text{{w}}_{\Delta i,\Delta j}=\delta_{\Delta i,\pm 1}\delta_{\Delta j,\pm 1}$) serves as a measure analogous to the RQA quantifier DET. It quantifies the prevalence of diagonal lines in an RP, which are characteristic of deterministic dynamics. Consequently, its value is negligible for Gaussian White Noise (GWN) but increases substantially for systems with memory, peaking for the deterministic chaotic and periodic systems (Logistic Map, Lorenz’63, and Sine function).
@@
Sides Motif: The “sides” motif ($\text{{w}}_{\Delta i,\Delta j}=\delta_{\Delta i,0}\delta_{\Delta j,\pm 1}+\delta_{\Delta i,\pm 1}\delta_{\Delta j,0}$) probes for vertical and horizontal lines, which correspond to laminar states or persistence in a region of phase space. This correlation is most pronounced for the strongly correlated AR(1) process and the chaotic systems. This indicates that in these systems, states have a tendency to recur in the same neighborhood for short periods, a direct consequence of a short-term linear memory.
@@
Anti-Diagonal Motif: The anti-diagonal motif ($\text{{w}}_{\Delta i,\Delta j}=\delta_{\Delta i,\pm 1}\delta_{\Delta j,\mp 1}$) provides the clearest distinction between system types. Such a pattern implies a form of time-reversal symmetry in recurrences ($x_{i}\approx x_{j+1}$ and $x_{i+1}\approx x_{j}$), a structure that is dynamically forbidden in autonomous, forward-evolving deterministic systems. Accordingly, the RPC value for this motif is nearly zero or slightly negative for all deterministic systems analyzed. In contrast, the stochastic AR(1) process exhibits a small but positive correlation. This does not imply a specific dynamical rule but rather that random fluctuations can coincidentally produce such patterns, whereas deterministic rules actively suppress them.
@@
The $\ell\text{RPC}$ method effectively dissects the mixed phase space of the Standard Map by acting as a dynamical filter (Fig. 6). We focus on a weight matrix with point-wise non-zero entries that takes the form $\text{{w}}_{\Delta i,\Delta j}=\delta_{\Delta i,k}\delta_{\Delta j,l}$ for various lags $k$ and $l$. This pattern probes the correlation between a recurrence at $(i,j)$ and another recurrence at $(i+k,j+l)$.
Figure 6: Local Recurrence Pattern Correlation ($\ell\text{RPC}$) in the Standard Map phase space. The $\ell\text{RPC}$ is calculated for uniformly distributed initial conditions $(x_{0},y_{0})\in[0,2\pi]$ across the phase space for $K=2.5$ and, with RPs built using a fixed $\varepsilon=0.5$. Each panel shows the $\ell\text{RPC}$ for a different correlation motif defined by the weight matrix $\text{{w}}_{\Delta i,\Delta j}=\delta_{\Delta i,k}\delta_{\Delta j,l}$, where the time lags $(k,l)$ are indicated in the panel titles. Warm colors (red) denote high positive correlation, indicating that the chosen recurrence pattern is prevalent, while cool colors (blue) denote low or negative correlation – the chosen pattern is rare.
@@
Figure 7: Detection of inherent recurrence times using $\ell\text{RPC}$ in the Lorenz ’63 system with a sampling step of $dt=0.02$. Panel (a) shows the global RPC as a function of time lag $\Delta t$, alongside the standard autocorrelation function for comparison. The bottom panels depict the localized $\ell\text{RPC}$ over the Lorenz attractor, evaluated at time lags corresponding to prominent peaks in the RPC. Some peak lags coincide with the durations of known heteroclinic unstable periodic orbits (UPOs), indicated by vertical lines colored by period [28]. We highlight three UPOs with the shortest durations: (d) period-2 (LR), (f) period-3 (LLR), and (g) period-4 (LLLR). Panels (b), (c), and (e) illustrate $\ell\text{RPC}$ for fractional durations of UPOs, which also align with secondary RPC peaks, and are related to orbits trapped in a single wing. The recurrence plot was constructed with a threshold such that only $1\%$ of the trajectory points are recurrent on average.
@@
Correlated recurrences are shown here to be directly associated with the topological organization of chaos and specific regions of phase space. This suggests that recurrent nonlinear dynamics are more effectively characterized by local properties rather than global statistics. Accordingly, time-indexed recurrence pattern quantifiers provide a more suitable and sensitive measure for capturing the underlying structure of nonlinear dynamical systems.
@@ IV Conclusions
Our primary contribution is to show that recurrence patterns encode fundamental properties of the underlying dynamical system beyond traditional recurrence-based quantifiers. The method’s adaptive sensitivity, which emphasizes structure in sparse RPs and non-recurrence in dense RPs, allows for a robust analysis across different dynamical regimes
@@
The principles of RPC can be extended beyond single time series to multivariate systems. Applying this framework to Cross Recurrence Plots would enable the geometric detection of structured, time-lagged interrelationships, providing a novel approach to quantify directional influence. This could reveal complex coupling patterns, such as identifying if a state in one system consistently precedes a state in another for distinct time lags, thereby uncovering the rules of their nonlinear interaction from time series measurements.

### v2

@@ Abstract
Recurrence plots (RPs) are powerful tools for visualizing time series dynamics; however, traditional Recurrence Quantification Analysis (RQA) often relies on global metrics, such as line counting, that can overlook system-specific, localized structures. To address this, we introduce Recurrence Pattern Correlation (RPC), a quantifier inspired by spatial statistics that bridges the gap between qualitative RP inspection and quantitative analysis. RPC is designed to measure the correlation degree of an RP to patterns of arbitrary shape and scale. By choosing patterns with specific time lags, we visualize the unstable manifolds of periodic orbits within the Logistic map bifurcation diagram, dissect the mixed phase space of the Standard map, and track the unstable periodic orbits of the Lorenz ’63 system’s 3-dimensional phase space. This framework reveals how long-range correlations in recurrence patterns encode the underlying properties of nonlinear dynamics and provides a more flexible tool to analyze pattern formation in recurrent dynamical systems.
@@ II Methods
is the average over all entries, known as the recurrence rate, so the denominator is the variance of the RP matrix. The weights $\text{{w}}_{\Delta i,\Delta j}$ are defined over time lags $\Delta i=i^{\prime}-i$ and $\Delta j=j^{\prime}-j$, so the correlation structure is translation invariant over the entire RP, making RPC a global measure of such generic motif. Moreover, the quantifier is normalized given the factor
@@
Assuming that the relatedness of time–series values is local, we restrict the lags to a finite window so that $\mathbf{w}_{\Delta i,\Delta j}$ is non-zero only within a bounded support. This prevents boundary artifacts and keeps the method computationally feasible even for larger weight matrices. Consequently, the inner summation is evaluated only over indices where the weights are non-zero. Near the boundaries of the RP, no issues arise because the normalization factor $W$ is computed only over the admissible combinations of time lags that fit inside the RP; outside this range, the corresponding weights are identically zero.
@@
In this context, we aim to develop a quantifier that captures not only global recurrence patterns but also those localized to individual values in the time series, and so in the state space. The local version can be conceptualized as applying the quantifier to individual columns of the RP, each associated with a specific time index $i$ in Eq. 2. Under a small recurrence-threshold condition, this correspondence can be approximated by a specific location in state space.
The local RPC (${\ell}\text{RPC}$) is defined as follows:
@@
As a result, the global RPC is not simply the average of the local values. This distinction makes the global definition better suited for identifying properties shared across the entire time series. The local RPC captures context-sensitive features by adapting to the local recurrence rate.
@@ III Results and Discussion
For one-dimensional time series, we applied the PECUZAL algorithm [25] for phase space reconstruction. This optimizes embedding parameters using non-uniform delays and a cost function tailored to the embedding dimension, without relying on a threshold parameter.
@@
The simplest correlated patterns arise from combinations of nearby recurrences, defined by the weight matrix $\text{{w}}_{\Delta i,\Delta j}=1$ only when $\Delta i=\pm 1$ or $\Delta j=\pm 1$. This corresponds to single time-lag recurrences aligned with vertical, diagonal, or anti-diagonal directions. Such a definition is commonly used in spatial correlation analysis, where long-range temporal causality is disregarded in favor of identifying nearby values in spatial clusters. We apply this simple weight matrix to representative dynamical systems [Fig. 3(a)], revealing distinct recurrence fingerprints that differentiate stochastic from deterministic dynamics.
@@
Figure 3: Recurrence Patterns Correlation (RPC) for paradigmatic dynamics and (a) representative recurrence motifs, and (b) as a function of the time series length. Here we consider a $1\%$ recurrence rate across all recurrence plots. In this scale, the recurring condition is highlighted over nonrecurring patterns. Each bar represents a specific weight matrix $\text{{w}}_{\Delta i,\Delta j}$, corresponding to: sides ($\text{{w}}_{\Delta i,\Delta j}=\delta_{\Delta i,0}\delta_{\Delta j,\pm 1}+\delta_{\Delta i,\pm 1}\delta_{\Delta j,0}$), diagonals ($\text{{w}}_{\Delta i,\Delta j}=\delta_{\Delta i,\pm 1}\delta_{\Delta j,\pm 1}$), and anti-diagonals ($\text{{w}}_{\Delta i,\Delta j}=\delta_{\Delta i,\pm 1}\delta_{\Delta j,\mp 1}$). In the bottom panel, the solid line represents the mean over 20 trials, and the error bars indicate the corresponding standard deviation. In the top panel, we present results for time series with length $N=10000$.
We investigated how the RPC quantifier depends on the amount of available data by computing it for increasing time series lengths and averaging over 20 independent trials [Fig. 3(b)]. For sufficiently long time series, the RPC converges to a stable value with low standard deviation, whereas short series exhibit significant variability due to insufficient sampling. This analysis demonstrates that the method requires a minimal data length to obtain statistically reliable estimates and clarifies the conditions under which the measure approaches its asymptotic value.
@@
Diagonal Motif. The correlation for the diagonal motif ($\text{{w}}_{\Delta i,\Delta j}=\delta_{\Delta i,\pm 1}\delta_{\Delta j,\pm 1}$) serves as a measure analogous to the RQA quantifier DET. It quantifies the prevalence of diagonal lines in an RP, which are characteristic of deterministic dynamics. Consequently, its value is negligible for Gaussian White Noise (GWN) but increases substantially for systems with memory, peaking for the deterministic chaotic and periodic systems (Logistic Map, Lorenz’63, and Sine function).
@@
Sides Motif. The “sides” motif ($\text{{w}}_{\Delta i,\Delta j}=\delta_{\Delta i,0}\delta_{\Delta j,\pm 1}+\delta_{\Delta i,\pm 1}\delta_{\Delta j,0}$) probes for vertical and horizontal lines, which correspond to laminar states or persistence in a region of phase space. This correlation is most pronounced for the strongly correlated AR(1) process and the chaotic systems. This indicates that in these systems, states have a tendency to recur in the same neighborhood for short periods, a direct consequence of a short-term linear memory.
@@
Anti-Diagonal Motif. The anti-diagonal motif ($\text{{w}}_{\Delta i,\Delta j}=\delta_{\Delta i,\pm 1}\delta_{\Delta j,\mp 1}$) provides the clearest distinction between system types. Such a pattern implies a form of time-reversal symmetry in recurrences ($x_{i}\approx x_{j+1}$ and $x_{i+1}\approx x_{j}$), a structure that is dynamically forbidden in autonomous, forward-evolving deterministic systems. Accordingly, the RPC value for this motif is nearly zero or slightly negative for all deterministic systems analyzed. In contrast, the stochastic AR(1) process exhibits a small but positive correlation. This does not imply a specific dynamical rule but rather that random fluctuations can coincidentally produce such patterns, whereas deterministic rules actively suppress them.
@@
We note that because our method relies on RPs, detecting a specific type of dynamics requires that (i) the system changes its states during its evolution - so that the RP contains informative structures - and (ii) the trajectory visits the phase-space region where the corresponding property is present. This explains why the Logistic Map exhibits a low $\ell\text{RPC}$ in the period-one regime ($r<3.0$) and, similarly, why the RPC does not detect a period-one UPO when the system is currently in a higher-period periodic orbit.
@@
The $\ell\text{RPC}$ method effectively dissects the mixed phase space of the Standard Map by acting as a dynamical filter (Fig. 6). We focus on a weight matrix with point-wise non-zero entries that takes the form $\text{{w}}_{\Delta i,\Delta j}=\delta_{\Delta i,k}\delta_{\Delta j,l}$ for various lags $k$ and $l$ combinations. This pattern probes the correlation between a recurrence at $(i,j)$ and another recurrence at $(i+k,j+l)$.
Figure 6: Local Recurrence Pattern Correlation ($\ell\text{RPC}$) in the Standard Map phase space, considering single time-lag motifs. The $\ell\text{RPC}$ is calculated for uniformly distributed initial conditions $(x_{0},y_{0})\in[0,2\pi]$ across the phase space for $K=2.5$ and, with RPs built using a fixed $\varepsilon=0.5$. Each panel shows the $\ell\text{RPC}$ for a different correlation motif defined by the weight matrix $\text{{w}}_{\Delta i,\Delta j}=\delta_{\Delta i,k}\delta_{\Delta j,l}$, where the time lags $(k,l)$ are indicated in the panel titles. Warm colors (red) denote high positive correlation, indicating that the chosen recurrence pattern is prevalent, while cool colors (blue) denote low or negative correlation – the chosen pattern is rare.
We display each panel in Fig. 6 so that its position matches the time–lag tuple $(\Delta i,\Delta j)$ it represents in the recurrence plot. In this arrangement, motifs associated with diagonal patterns (i.e., equal time lags) are aligned along the main diagonal of the figure, while motifs associated with vertical patterns are aligned vertically. This layout provides an intuitive visual mapping between a subplot’s position and the recurrence structure it encodes.
@@
Figure 7: Detection of inherent recurrence times using $\ell\text{RPC}$ in the Lorenz ’63 system with a sampling step of $dt=0.02$. Panel (a) shows the global RPC as a function of time lag $\Delta t$, alongside the component $x$ autocorrelation function for comparison. The bottom panels depict the localized $\ell\text{RPC}$ over the Lorenz attractor, evaluated at time lags corresponding to prominent peaks in the RPC. Some peak lags coincide with the durations of known heteroclinic unstable periodic orbits (UPOs), indicated by vertical lines colored by period [28]. We highlight three UPOs with the shortest durations: (d) period-2 (LR), (f) period-3 (LLR), and (g) period-4 (LLLR). Panels (b), (c), and (e) illustrate $\ell\text{RPC}$ for fractional durations of UPOs, which also align with secondary RPC peaks, and are related to orbits trapped in a single wing. The recurrence plot was constructed with a threshold such that only $1\%$ of the trajectory points are recurrent on average.
@@
Correlated recurrences are shown here to be directly tied to the topological organization of chaos and to specific regions of phase space. This indicates that recurrent nonlinear dynamics are more effectively characterized through local properties rather than global statistics. Time-indexed recurrence pattern quantifiers therefore provide a complementary and sensitive perspective for revealing the fine-scale geometric structures embedded in recurrence plots.
Our results indicate that different dynamical regimes generate distinct recurrence motifs, implying that the weight matrix acts as a tunable lens for emphasizing system-specific geometry. Rather than selecting motifs manually, an important extension is to identify weight matrices that enhance contrast for a given system or highlight particular dynamical features. Preliminary evidence suggests that simple searches over candidate motifs already recover structures consistent with UPO periods or invariant-manifold geometry, motivating a data-driven strategy for inferring the intrinsic recurrence organization. Because $\ell\text{RPC}$ is highly sensitive to the spatial arrangement of recurrence points, it supports broader automated exploration of motif space without prescribing matrix forms in advance. Such inferred motifs would complement traditional line-based measures, extending the range of detectable structures arising from nonlinear evolution.
@@ IV Conclusions
Our primary contribution is to demonstrate that recurrence patterns encode fundamental properties of the underlying dynamical system that extend beyond traditional recurrence-based quantifiers. The method’s adaptive sensitivity, which emphasizes structure in sparse RPs and non-recurrence in dense RPs, allows for a robust analysis across different dynamical regimes
@@
Future work should focus on developing agnostic strategies to infer weight matrices that represent the system’s preferred recurrence motifs, rather than selecting them a priori. Combined with multivariate and cross-recurrence formulations, this would extend RPC toward detecting structured, time-lagged interactions across variables, enabling a compact geometric characterization of directional influence in complex nonlinear systems.